# AI paintings vs. Human Paintings? Deciphering Public Interactions and Perceptions towards AI-Generated Paintings on TikTok

### v1

@@ Introduction
Video quality is an important control variable in this paper, and its impact on user engagement has been confirmed by many previous studies(Diallo, Fieau, and Hennequin 2014). So, in the process of sample preprocessing, we also considered the video quality and the flowchart is presented in Fig. 2b. Since image-text videos only contain Paintings and text, all videos have the same page layout and presentation method, so here, video quality is equivalent to the aesthetic quality of the Paintings in the videos. Firstly, we downloaded all the Paintings from the two types of videos, removed duplicate Paintings within the same video, and obtained 2383 AI paintings and 1153 human paintings. In our research, we used the improved-aesthetic-predictor (LAION-Aesthetics V2) [^5] https://laion.ai/blog/laion-aesthetics/ as the IAQA model. The model’s training process is divided into two steps: embedding the input images using Clip ViT/14 (Radford et al. 2021) and then using MLP to train the image data from three well-known datasets, namely, 176,000 image-text pairs from SAC (Pressman, Crowson, and Contributors 2023) dataset, 15,000 image-text pairs LAION-Logos dataset with aesthetic scores from 1 to 10, and 250,000 photos from the AVA dataset with aesthetic scores from 1 to 10. The data after aesthetic quality scoring of 2.73 billion images in the LAION 5B dataset (Schuhmann et al. 2022) by this Predictor trained Stable Diffusion V1666https://github.com/CompVis/stable-diffusion/tree/ce05de28194041e030ccfc70c635fe3707cdfc30#stable-diffusion-v1 , one of the most professional AIGC models currently. Therefore, this model’s comprehensive reliability and accuracy in assessing image aesthetic quality are likely the best now available.
To control the quality of the two types of videos at the same level and avoid the interference of extreme values, we excluded videos with less than $4.5andmorethan$6.5 average image aesthetic quality scores, ultimately obtaining 207 AIGP videos and 210 human painting videos. The independent sample T-test results (Table 3 in Appendices) also confirmed that there was no significant difference in IAQA values between the two types of videos (t=1.499, P = $0.135>$0.05). We then divided the videos into four quality intervals: Q1: (4.5,5.0], Q2: (5.0,5.5], Q3: (5.5,6.0] and Q4: (6.0, 6.5]. We illustrate some Paintings from different aesthetic intervals in Fig.3. By counting the total number of views of the two types of videos, we find that their total number of views has reached hundreds of millions, which indicates that a large number of users are viewing similar videos on TikTok. Therefore, our sample data is universal and reliable.
@@
For topic modeling, we initially filtered the comments on AIGP videos to extract all negative comments, resulting in 19,071 comments. These comments were subjected to topic modeling in this study. After comparing several mainstream topic modeling methods, such as Latent Dirichlet Allocation (LDA), Latent Semantic Analysis (LSA), and Non-Negative Matrix Factorization (NMF), we opted for the NMF model because our data contained amount of short texts but not long documents (Gamage et al. 2022). We trained the NMF model for different numbers of topics from k = 1 to 15, and we calculated the mean topic coherence across all topics to determine the optimal number of topics automatically (Figure. 6 in Appendices). Semantic coherence involves examining the co-occurrence of words within a document to ensure that the topics generated by the model possess semantic cohesion  (Wang and Lu 2023), a mainstream method for assessing topic quality. When k = 1, semantic coherence was the highest, but a single topic could not accurately summarize our collected comment data. Hence, we chose 14 topics for the final training phase, as this number demonstrated a relatively high level of semantic coherence.
Results
Table 1: The user engagement calculated by traditional formula. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and Wilcoxon test otherwise.
Measures Normality Test types Items Mean Std W p
User engagement N Wilcoxon test AIGP videos (N = 207) 0.12 0.07 -13.774 0.00***
HP videos (N = 210) 0.23 0.08
Measures Normality Test types Items Mean Std W p
@@ Results
Informed by our hypotheses, all tests in the study were conducted using one-tailed comparisons. The analysis of our data and manipulation checks utilized independent T-tests, Wilcoxon tests, or one-way ANOVA, depending on whether the data distribution was normal. Our normality tests on user engagement for both types of videos revealed that the data did not conform to a normal distribution. As a result, we employed the Wilcoxon test to compare the user engagement of the two sample types, as detailed in Table 1  (Hu et al. 2022). Table 2 compares the comment sentiment of two types of videos. The results supported hypotheses H1 and H3, indicating that people’s user engagement (W = -13.774, P = 0.00) and sentiment scores for comments (W = -36.725, P = 0.00) on human painting videos were significantly higher than those for AIGP videos. These findings address RQ1 and RQ3-part1 posed in the paper.
Regarding the impact of video aesthetic quality on user engagement, the videos’ average aesthetic quality was divided into high, medium, and low ranges according to the LAION 5B  (Schuhmann et al. 2022) divisions for IAQA: (5.5, 6.5], (5.0, 5.5], and (4.5, 5.0]. We then assessed the impact of these quality ranges on user engagement, as detailed in Table 4 in Appendices. As this involved comparisons among three samples, we conducted one-way ANOVA tests on user engagement corresponding to different quality intervals for AIGP videos and human drawing videos, respectively. The results indicated that image aesthetic quality does not significantly affect user engagement (F = 2.056, P = 0.131; F = 1.663, P = 0.192). Moreover, we found that the difference in user engagement between the two types of videos in different quality intervals is also very close to each other. The difference in high IAQ is 0.123, in middle IAQ is 0.122, and in low IAQ is 0.133. Hence, the experimental results indicate that hypothesis H2 is supported, and RQ2 posed in the paper is answered.
RQ1: How are people’s intentions to engage with AIGP on social media platforms?
Stronger Desire to Share AIGP
We calculated the like, comment, favorite, and share ratios of each video based on views in Table 5. We found that the like, comment, and favorite ratios for human painting videos significantly exceed those for AIGP videos, particularly in the comment ratios of 0.001 (AIGP) and 0.003 (HP). These data corroborate our previous results on user engagement comparison, and also show that using machine learning methods to calculate user engagement is reliable. Interestingly, the share ratio of AIGP videos is much higher than that of human painting videos. Based on our speculation, the possible reason is that the ability of AI painting breaks people’s perception of AI technology, and the speed of its development is also shocking to people who want to share it and discuss it with their friends. However, as with comments, we are not clear about users’ motivations for sharing, which can be positive signals or negative. Therefore, we analyzed the comment motivation specifically in a later section.
We compared the impact of different aesthetic quality intervals of Paintings on user engagement for both video types, and the final results found no significant correlation between the two. This suggests that when browsing different artworks on social media, people may be more concerned with the intrinsic meaning and value of the paintings than their glossy appearance. Moreover, we noticed that the engagement with human painting videos increases slightly as aesthetic quality decreases. Observing the comments on human painting videos, there appears to be a lot of positive social support for the creators, such as when amateur painters upload less-than-perfect paintings and drawings on TikTok. Interestingly, many users offer encouragement rather than criticism. Some examples are as follows:
The Drawing Looks Better Than The Original!
your art just gets better and better I’m amazed!
RQ3: How are public sentiments toward AIGP? What motivates people’s negative attitudes toward AIGP?
After comparing the comment sentiment scores of both video types, we found in Table 2 that human painting videos have significantly more positive comments than AIGP videos. It is undeniable that the number of positive comments about both types of paintings is higher than the number of negative comments, and this data suggests, to a certain extent, that many people have already accepted that AI painting exists. Some AI-created artworks are also well-liked. However, compared to human artworks, some people tend to harbor more negative emotions when discussing AIGP content on social media. Overall, people still prefer to see human artworks on social media platforms rather than AI-generated ones.
This section addresses RQ3-part 2 by identifying seven types of negative comments and reasons for reluctance to interact with or express negative sentiments toward AIGPs on social media. The topic modeling revealed 14 common topics within the dataset, among which the top 7 topics with the highest proportion are listed in Table 6. To avoid the influence of positive comments, all topics were extracted from 19,071 negative comments about AIGP posted on TikTok.
It is too real/appalling to be unacceptable
After analyzing the topics, we found a slight overlap between Topic 1 and Topic 7, and we discussed these two topics together. These topics reflected that the reason for people’s negative attitude toward AIGP is that AIGP has become too realistic and is perceived likely to affect our lives in the future. Indeed, the advancement of AI technology has become increasingly unpredictable, and we are facing a dilemma that it is difficult to distinguish between human paintings and AIGPs. This dilemma will become increasingly challenging to break free from, along with the development of AI technology. One controversial incident involved using a deepfake voice in the documentary film “Roadrunner” about celebrity chef Anthony Bourdain (Gamage et al. 2022). The film incorporated an artificially artificial voice to simulate Bourdain speaking, which he never originated or consented to. This means that we may not even know when others have privately used our voices and likenesses. So when users are browsing these AIGPs, it’s hard for them not to conjure up Paintings of the day when we can no longer tell which ones are AI-generated and which ones are human’s own. Examples are given below:
If you don’t say AI, you won’t know AI.
Finally, it’s getting harder for AI to discriminate…
I feel like I’m in a spiral of self-contradictions
Topic 2 expresses the idea that when many users see AIGPs, they are indeed attracted by the fine character portrayal, rich color palette, and other artistic features, but when they see the AI tags underneath the video, they start to hesitate and sway from side to side or even feel disgusted and ashamed of the paintings, like these typical comments:
This looks so gorgeous! Shame it’s AI Art tho
the way my smile went away when i noticed its AI
I am not acceptable for AI paintings, and it’s not art.
Sadly AI art isn’t art, this is just slide after slide of dogshit
if it’s made by AI it’s not art
I hate ai soo much
The fact that its AI generated makes it boring and empty
It is not good enough to meet my expectations.
Many AIGPs are generated based on real people, anime characters, or classic scenes. The concept of Topic 4 is that many people think that AI-generated characters in Paintings don’t live up to their expectations or don’t capture some of the character’s traits, which leads to their negative attitudes about AIGP. Some comments are given below:
i think AI did kyo n yuki dirty nooo
AI inconsistency when it comes to skin colour
It is so scary and weird.
There are two meanings we can get from Topic 5. Firstly, the word “dark” indicates that there may be more dark and eerie types of AIGPs on TikTok, and this style of drawing makes people feel uncomfortable. On the other hand, AI-generated technology has always had some complex technical difficulties, such as the drawing of human limbs and skin color:
Some of the keywords in this topic, like “potter” and “hand” indicate that some human-themed AIGPs on social media have weird limbs or postures. These Paintings are likely to horrify and disturb users, leading to negative comments.
It’s plagiarism
The reason for the negative attitude towards AIGP, as explained in Topic 6, is that AIGP’s creative process is perceived as a form of plagiarism, and therefore, people feel cheated. Some typical examples as follows:
AI art is theft
People think that why AI can generate high-quality paintings because of plagiarising the artists’ paintings, which are not thought out by themselves, and therefore, AI painting should not be accepted and spread.
Discussion
Hesitation and Tentativeness in Interaction with AIGP
According to Nadarzynski’s study of the public response to AI-led chatbot services, “AI hesitancy” was prevalent during users’ interaction with chatbots  (Nadarzynski et al. 2019). It demonstrated that a substantial number of people may feel hesitant to use AI modules, considering the accuracy and security of these services. By comparing the degree of users’ interaction between the AIGP video and the human painting video, we found that people still feel dithery when responding to the AIGP on social media, although the interaction does not involve services or device usage. This low level of willingness to engage is reflected in the likes, favorites, and comments on AIGP-type videos. Taking into account possible existing technical reasons for generative AI, we also controlled the moderating variable of IAQ. However, the moderating role of IAQ is not significant. This finding reinforces the fact that there may still be an unacceptable attitude and mistrust in regard to the generative AI technology itself. Introducing the Fear theory into the AI-related research, several studies have acknowledged that new technology is frequently perceived as a potential threat due to the risks it may pose to individuals or society as a whole. (Osiceanu 2015; Khasawneh 2018). Regarding AI specifically, we are currently undergoing a shift towards cities and societies that are governed by various artificial intelligences (Cugurullo and Acheampong 2023). However, the overall risks of this transition are still not fully understood (Yigitcanlar et al. 2020). Based on these previous unresolved questions, this paper contributes interesting empirical findings to research related to generative AI risks and social attitudes. We derived seven topics from users’ comments on AIGP to uncover why negative attitudes still exist. Among the topics, a negative attitude due to the fear of AIGP was confirmed in this study. We also find that ambivalence and looking too real were typical reasons that led to negative attitudes.
Influence of AIGP to Art
There is no doubt that AIGP has had a significant impact on the art industry, as can be seen from the results of topic modeling, with Topic 1 and Topic 7 all highlighting the fact that today’s AIGPs are so similar to human paintings that we cannot distinguish some of the high-quality AIGPs from human paintings. On December 6, 2023, a post made it to the top of Weibo (one of China’s most famous social media platforms, similar to Twitter) hotlist, which was along the lines of the mascot for the Spring Festival Gala announced by China Central Broadcasting Television (CCTV) being thought to have been generated by AI. Some users even posted long posts analyzing various details to prove that it was generated by AI. Although an official announcement was made to reply to the challenge and the designer’s original design was released, users still firmly believed it was AI-generated and accused of this behavior. As a result, the emergence of AIGP has added many burdens to artists. As far as painters are concerned, they not only have to design creative and high-quality artworks but also need to worry whether their works will be labeled as AI and blamed. Especially, some high-level painters have absolute confidence that their paintings can be better than AIGPs. Still, they are afraid that the paintings they spent a lot of time and energy to create will be regarded as AI-generated. Therefore, this may be one of the reasons why AIGP is likely to cause many excellent artists to lose their creative drive and support anti-AI events.
Implications for social media platform design
In addition, we analyzed and compared the sentiment of comments on AIGP and human paintings. The results showed that positive comments generally dominated people’s comments on both types of paintings, but the rate of negative comments on AIGP videos was significantly higher than that on videos of human paintings. It means that there are still a lot of people who have a negative attitude towards AIGP, and many people are likely unwilling to see AI-generated Paintings on social media platforms. Therefore, in order to further explore what makes people hold negative attitudes towards AIGP content on TikTok, we conducted topic modeling of negative AIGP comments and finally identified seven main reasons. These reasons provide suggestions and guidance for optimizing future AI-generated models, from which we can analyze people’s preferences to determine which AIGP to put on social media to attract users’ attention and promote an intimate connection between the platform and users.
Limitations and Future Work
This study uses user-generated content (UGC) on the TikTok platform as the source of experimental data, so data collection was done on a single social media platform. However, whether there is a difference in the perception of AIGP among users of different platforms, such as Twitter, Facebook, Reddit, etc., is an area not considered in this paper. However, it is undeniable that the number of active users on TikTok is still the highest of all social media platforms, and all AIGP videos have to be tagged with “ai”, which very few platforms do. We also tried to spread our survey to other mainstream social media. Still, we found that many platforms are not open and transparent about their data, such as the number of video videos or post views, which is an essential measure of user engagement in this experiment. Thus, this can cause significant difficulties in collecting data for our experiment, and the accuracy of the data cannot be guaranteed. At the same time, we did not focus only on numerical data on the platform. We also processed and analyzed tens of thousands of comments, which reflect the actual attitudes and emotions of platform users towards the two types of paintings. Consequently, integrating these several methods and data can guarantee the rigor of the experimental results. To further validate our findings, our future work will no longer be limited to social media platforms, and we hope to investigate and analyze these issues in this paper in more detail through questionnaires or real-life interviews.
Conclusion
Because of TikTok’s huge user groups and development speed far exceeding that of other platforms such as Facebook and Twitter, people’s interactions and perceptions towards AIGP on the TikTok platform can reflect a social phenomenon to a certain extent. Therefore, this paper investigates the user engagement and emotional tendency of comments on AIGP on social media based on the TikTok platform and compares it with human drawings. Finally, it is concluded that people are more willing to interact with the content of human paintings on social media and have higher positivity towards their comments. We also considered the potential moderating role of the IAQ in human-AIGP interactions, and the results did not support this hypothesis. Finally, seven main themes derived from the topic modeling, including looks too real, ambivalence, etc., indicated why some people still have negative perceptions towards AIGP.
Appendix A Appendices
Measures Normality Test types Items Mean Std Video views t p
IAQA Y Independent T-test AI videos (N = 207) 5.48 0.46 259088700 1.499 0.135
HP videos (N = 210) 5.41 0.48 40870900
Table 3: A comparison of IAQA between AIGP videos and human painting (HP) videos. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and the Wilcoxon test otherwise.
Measures Test types IAQA AIGP videos (N = 207) Human painting videos (N = 210)
Sample
size
Mean std F P
Sample
size
Mean std F P Difference
User engagement (Formula)
One-way
ANOVA
(5.5,6.5] 109 0.13 0.068 1.788 0.17 84 0.23 0.072 0.016 0.984 0.1
One-way
ANOVA
(5.0,5.5] 66 0.11 0.069 80 0.22 0.081 0.11
(4.5,5.0] 32 0.12 0.069 46 0.22 0.09 0.1
Table 4: A comparison of the user engagement from different video quality intervals. We used the One-way ANOVA test for three samples.

### v2

@@ Introduction
Video quality is an important control variable in this paper, and its impact on user engagement has been confirmed by many previous studies(Diallo, Fieau, and Hennequin 2014). So, in the process of sample preprocessing, we also considered the video quality. Since image-text videos only contain Paintings and text, all videos have the same page layout and presentation method, so here, video quality is equivalent to the aesthetic quality of the Paintings in the videos. Firstly, we downloaded all the Paintings from the two types of videos, removed duplicate Paintings within the same video, and obtained 2383 AI paintings and 1153 human paintings. In our research, we used the improved-aesthetic-predictor (LAION-Aesthetics V2) [^5] https://laion.ai/blog/laion-aesthetics/ as the IAQA model. The model’s training process is divided into two steps: embedding the input images using Clip ViT/14 (Radford et al. 2021) and then using MLP to train the image data from three well-known datasets, namely, 176,000 image-text pairs from SAC (Pressman, Crowson, and Contributors 2023) dataset, 15,000 image-text pairs LAION-Logos dataset with aesthetic scores from 1 to 10, and 250,000 photos from the AVA dataset with aesthetic scores from 1 to 10. The data after aesthetic quality scoring of 2.73 billion images in the LAION 5B dataset (Schuhmann et al. 2022) by this Predictor trained Stable Diffusion V1666https://github.com/CompVis/stable-diffusion/tree/ce05de28194041e030ccfc70c635fe3707cdfc30#stable-diffusion-v1, one of the most professional AIGC models currently. Therefore, this model’s comprehensive reliability and accuracy in assessing image aesthetic quality are likely the best now available.
Figure 4: Image Aesthetic Quality Schematic. This schematic illustrates images from both categories that were evaluated using the Image Aesthetic Quality Assessment (IAQA) tool, which assigned each image a quality score.
Variables Description
Dependent variable User Engagement (Ue) User Engagement refers to the level of user interaction with the content. In our study, we measure this variable using the number of comments, likes, and shares.
Independent variable Human Origin Level (Hum) Human origin level is a binary variable indicating the origin of a given painting. It is coded as 0 for AIGP and 1 for HP
Moderator Aesthetic Quality (Aes) Image Aesthetic Quality represents the aesthetic quality of the painting. We measure this variable using aesthetic quality scores.
Table 1: Descriptions of variables in the regression model
After collecting and measuring the IAQ of approximately 3,500 images of AIGP and HP published on the TikTok platform, we found that the IAQ scores of both types of paintings predominantly ranged between 4.5 and 6.5. To ensure comparability and avoid the influence of extreme values on the experimental results, we excluded videos with an average image aesthetic quality score lower than 4.5 or higher than 6.5. This decision was made for two main reasons: (1) extremely low or high scores often corresponded to outlier cases—such as distorted, poorly generated, or overly edited images—which may introduce noise and distort the analysis; and (2) the majority of both AIGP and HP samples were concentrated within this score range, making it a representative and comparable interval for analysis. After applying this filtering criterion, we obtained 207 AIGP videos and 210 HP videos for further analysis. We illustrate some Paintings from different aesthetic quality intervals in Fig.4. By counting the total number of views of the two types of videos, we find that their total number of views has reached hundreds of millions, which indicates that a large number of users are viewing similar videos on TikTok. Therefore, our sample data is universal and reliable.
To test H2, we built an OLS regression model shown below and used Stata to examine whether IAQ moderates the relationship between user engagement and painting sources. Table 1 shows the descriptions of each variable in the regression model. The dependent variable is User Engagement (Ue), which captures the level of user interaction with each video. This variable is operationalized using observable engagement metrics such as the number of likes, comments, and shares. The independent variable is Human Origin Level (Hum), which distinguishes between two types of visual content: AIGP as 0 and HP as 1. This variable allows us to examine differences in engagement based on the origin of the artwork. In addition, Image Aesthetic (Aes) is introduced as a moderator variable, reflecting the aesthetic quality of each image. This variable is measured using quantitative aesthetic quality scores and is used to test whether aesthetic appeal moderates the relationship between painting source and user engagement. These variables collectively form the basis for exploring how the source and aesthetic quality of artwork jointly influence user behavior on social media platforms.
Since all data cleaning and preprocessing procedures were already completed in the “Data Collection and Pre-processing” section, we do not repeat those steps here. However, to reduce potential multicollinearity and enhance the stability of the regression analysis, we applied mean-centering to the continuous moderator variable, Aes. Centering was performed by subtracting the sample mean from each observed value of the variable, as shown below:
$\text{Aes}_{\text{centered}}=\text{Aes}_{i}-\overline{\text{Aes}}$ (1)
Finally, we build the following OLS regression model and analyze it using state, as shown below:
$Ue=\beta_{0}+\beta_{1}\cdot\text{Hum}+\beta_{2}\cdot\text{Aes}_{\text{centered}}+\beta_{3}\cdot\text{Hum}\cdot\text{Aes}_{\text{centered}}+\varepsilon_{i}$ (2)
@@
Figure 5: Sentiment Analysis Flowchart. This flowchart demonstrates the process of performing sentiment analysis on comments from AIGP and HP. Initially, 392,812 comments from AIGI videos and 63,445 comments from human painting videos were collected. Comments containing only “@” were removed, and comments with at least one like were retained, resulting in 42,199 comments for AIGI videos and 29,763 comments for human painting videos. Sentiment analysis was conducted using an NLP model, where comments were classified as positive (1), neutral (0), or negative (-1). The average sentiment values for AIGI and human painting videos were calculated and compared.
Measures Normality Test types Items Mean SD t p
User engagement N Independent T-test AIGP videos (N = 207) 0.104 0.058 -15.064 0.00***
HP videos (N = 210) 0.198 0.068
Table 2: The user engagement calculated by traditional formula. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and the Wilcoxon test otherwise.
Measures Normality Test types Items Mean SD W p
@@
For topic modeling, we initially filtered the comments on AIGP videos to extract all negative comments, resulting in 19,071 comments. These comments were subjected to topic modeling in this study. After comparing several mainstream topic modeling methods, such as Latent Dirichlet Allocation (LDA), Latent Semantic Analysis (LSA), and Non-Negative Matrix Factorization (NMF), we opted for the NMF model because our data contained a number of short texts but not long documents (Gamage et al. 2022). We trained the NMF model for different numbers of topics from k = 1 to 15, and we calculated the mean topic coherence across all topics to determine the optimal number of topics automatically (Fig. 6). Semantic coherence involves examining the co-occurrence of words within a document to ensure that the topics generated by the model possess semantic cohesion  (Wang and Lu 2023), a mainstream method for assessing topic quality. When k = 1, semantic coherence was the highest, but a single topic could not accurately summarize our collected comment data. Hence, we chose 14 topics for the final training phase, as this number demonstrated a relatively high level of semantic coherence.
Figure 6: Mean Coherence of Comments. This figure shows the mean coherence scores of comment topics as a function of the number of topics.
Results
To compare user engagement levels between AIGP and HP on TikTok, we conducted an independent samples t-test. Prior to the analysis, the normality of the user engagement data was assessed using the Shapiro-Wilk test. As the data met the assumption of normality, an independent t-test was deemed appropriate. The results reveal a statistically significant difference in user engagement between AIGP videos ($M=0.104,Std=0.058$) and HP videos ($M=0.198,Std=0.068$), $t(415)=–15.064,p<0.001$ in Table 2. The mean engagement for human-created paintings is substantially higher than that for AI-generated content. This finding suggests that users interact significantly more with human-created paintings compared to AIGP on social media platforms. These results indicate that H1 is accepted.
To examine whether IAQ positively moderates the relationship between Human Origin Level (Hum) and User Engagement (Ue), we conducted a regression analysis using Stata. The results are presented in Table 4. The coefficient for Hum is positive and statistically significant ($\beta=0.094,p<0.001$), indicating that HP is associated with higher levels of user engagement compared to AIGP, holding other variables constant. The coefficient for Aes is positive but not statistically significant ($\beta=0.008,p=0.387$), suggesting that aesthetic quality alone does not have a significant direct effect on user engagement in this model. Crucially, the interaction term Hum × Aes has a negative and non-significant coefficient ($\beta=–0.007,p=0.620$), indicating that IAQ does not significantly moderate the relationship between Hum and Ue. In other words, the effect of Hum on user engagement remains consistent regardless of variations in aesthetic quality. The overall model explains approximately 35.5% of the variance in user engagement ($R^{2}=0.355;AdjustedR^{2}=0.350$), suggesting a moderate level of explanatory power. In summary, while the type of painting source significantly influences user engagement, there is no evidence from this model that IAQ moderates this relationship. Hence, the experimental results indicate that hypothesis H2 is rejected.
To compare user comments on AI-generated paintings (AIGP) versus human-made paintings (HP) on social media, we conducted semantic analysis, including sentiment score comparisons and topic modeling based on the collected comments. Table 3 presents the results of the sentiment analysis. As the normality assumption was violated (as assessed by the Shapiro-Wilk test), a Wilcoxon rank-sum test—the non-parametric equivalent of the independent t-test—was employed. The results indicate a statistically significant difference in sentiment scores between the two groups. Specifically, comments on AIGP videos ($N=207$) had a mean sentiment score of 0.30 ($SD=0.93$), while comments on HP videos ($N=210$) had a higher mean sentiment score of 0.53 ($SD=0.83$). The Wilcoxon test yielded a test statistic of $W=-36.725$ with a p-value less than 0.001, indicating a highly significant difference. These results suggest that comments on HP content tend to be significantly more positive than those on AIGP content. In other words, users express more favorable sentiment toward content featuring human-made paintings compared to AI-generated ones, highlighting potential perceptual or emotional biases in user responses. Based on the sentiment analysis, hypothesis H3 is rejected.
Secondly, the mean coherence scores and the results of topic modeling are presented in Figure 6 and Table 6. Figure 6 shows a non-linear trend in coherence scores across different numbers of topics. Although the highest coherence score occurs at two topics, this result is often considered unreliable due to its limited interpretability and the overly broad grouping of themes. After a sharp decline at three topics, the coherence scores gradually increase as the number of topics rises, reaching a local peak at 14 topics. Based on these findings, we selected the model with 14 topics for further qualitative analysis, as it offers a balance between coherence and interpretability. From this model, we identified the top seven topics with the highest proportions, which are presented in Table 6.
Variables Model
Hum 0.094***
Aes 0.008
Hum*Aes -0.007
Constant 0.104
R-squared 0.355
Adj R-squared 0.350
Table 4: Table 4: Regression results. The results shows that while Human Origin Level affects user engagement, IAQ does not moderate this relationship. Note: *p<0.05, **p<0.01, ***p<0.001